# GNN-DT: Graph Neural Network Enhanced Decision Transformer for Efficient Optimization in Dynamic Environments

### v1

@@ 4 GNN-based Decision Transformer
Figure 1: Overview of the GNN-DT architecture. The input sequence, comprising return-to-go, action, and state, is processed through specialized embedding modules. The action graph $\mathcal{G}^{\mathbf{a}}_{t}=(\mathcal{N}^{\mathbf{a}}_{t},\mathcal{E}^{\mathbf{a}}_{t})$, with nodes $\mathcal{N}^{\mathbf{a}}_{t}\subset\mathcal{N}_{t}$, and the state graph $\mathcal{G}_{t}=(\mathcal{N}_{t},\mathcal{E}_{t})$ are encoded using GNN-based embedders to produce embeddings of dimension $F_{L}$. These embeddings serve as inputs to a GPT-2–based causal transformer, which predicts the next action token. The predicted action token acts as a decoder, generating actions by multiplying with specific GNN state node embeddings. This architecture enables efficient decision-making in scenarios with dynamic states and actions.
@@ 4.1 Sequence Embeddings
In detail, the State Embedder consists of $L$ consecutive Graph Convolutional Network (GCN) (Kipf & Welling, 2016) layers, which aggregate information from neighboring nodes as follows:
@@
where $\boldsymbol{x}_{t}^{(l)}\in\mathbb{R}^{N_{t}\times F_{l}}$ denotes the node embeddings at layer $l$ with $N_{t}$ number of nodes, $W^{(l)}\in\mathbb{R}^{F_{l}\times F_{l+1}}$ are trainable weights, $\sigma(\cdot)$ is a nonlinear activation (ReLU), $A_{t}$ is the adjacency matrix of the state graph $\mathcal{G}_{t}$, and $D$ is the degree matrix for normalization. After the final layer, a mean-pooling operation produces a fixed-size state embedding:
$\boldsymbol{\widetilde{s}}_{t}=\frac{1}{\lvert\mathcal{N}_{t}\rvert}\sum_{n\in\mathcal{N}_{t}}\boldsymbol{x}_{n,t}^{(L)},$ (7)
where $\boldsymbol{x}_{n}^{(L)}$ is the embedding of node $n$ at the $L$-th layer. This pooling step ensures that the state embedding is invariant to the number of nodes in the graph, enabling the architecture to scale with any number of EVs or chargers.
Similarly, the Action Embedder processes the action graph $\mathcal{G}^{\mathbf{a}}_{t}=(\mathcal{N}^{\mathbf{a}}_{t},\mathcal{E}^{\mathbf{a}}_{t})$ through $C$ GCN layers followed by mean pooling, producing the action embedding $\widetilde{\textbf{a}}_{t}$. All embedding vectors (states, actions, or the return-to-go value) have the same dimensions. This design leverages the dynamic and invariant nature of GCN-based embeddings, allowing the DT to handle variable-sized graphs.
@@ 4.2 Decoding Actions
Once the embedding sequence of length $K$ is constructed [^3] During inference the action ($\mathbf{a}_{t}$) and RTG ($\boldsymbol{\widehat{R}}_{t}$) of the last step $t$ are filled with zeros as they are not known., it is passed through the causal transformer GPT-2 (Chen et al., 2021) to produce a fixed-size output vector $\boldsymbol{y}_{t}\in\mathbb{R}^{F_{L}}$ for each step. Because DT architectures inherently generate outputs of fixed dimensions, an additional mechanism is required to manage dynamic action spaces. To address this, GNN-DT implements a residual connection that merges the final GCN layer embeddings $\boldsymbol{x}_{t}^{(L)}$ with the transformer output $\boldsymbol{y}_{t}$ for every step of the sequence.
Specifically, for each node $n\in\mathcal{N}_{t}^{\mathbf{a}}$, we retrieve its corresponding state embedding $\boldsymbol{x}_{n,t}^{(L)}\in\mathbb{R}^{1\times F_{L}}$ and multiply it with the transformer output token $\boldsymbol{y}_{t}\in\mathbb{R}^{1\times F_{L}}$, yielding the final action for node $n$:
$\hat{\mathrm{a}}_{n,t}=\boldsymbol{y}_{t}^{\mathsf{T}}\cdot\boldsymbol{x}_{n,t}^{(L)}.$ (8)
By repeating Eq. 8 for every step $t$ and every node $n\in\mathcal{N}^{\mathbf{a}}_{t}$ the final action vector $\hat{\mathbf{a}}_{t}$ is generated. This design allows the model to maintain a fixed-size output from the DT while dynamically adapting to any number of nodes (and hence actions). It effectively combines the high-level context learned by the transformer with the node-specific state information captured by the GNN, enabling robust, scalable decision-making even as the graph structure changes.
Figure 2: Training performance comparison of DT variations (classic DT, Q-DT, and GNN-DT) across 3 different datasets with 10.000 trajectories each and classic online RL algorithms for the same problem.
4.3 Action Masking and Loss Function
In GNN-DT, the learning of infeasible actions, such as charging an unavailable EV, is avoided through action masking. At each time step $t$, a mask vector $\mathbf{m}_{t}$, which has the same dimension as $\mathbf{a}_{t}$, is generated with zeros marking invalid actions and ones marking valid actions. For example, an action is invalid when the $a_{i,t}\neq 0$ and no EV is connected at charger $i$. The mean squared error between the predicted actions $\widehat{\mathbf{a}}_{t}$ and ground-truth actions $\mathbf{a}_{t}$ from expert or offline trajectories is employed as the loss function. For a window of length $K$ ending at time $t$, training loss is defined as:
$\mathcal{L}=\frac{1}{K}\sum_{\tau=t-K}^{t}\|\left(\widehat{\mathbf{a}}_{\tau}-\mathbf{a}_{\tau}\right)\circ\mathbf{m}_{\tau}\|^{2}.$ (9)
By incorporating the mask into the loss calculation (elementwise multiplication), a focus solely on valid actions is enforced, thereby preserving meaningful gradient updates and enhancing training stability.
@@ 5 Experiments
In this section, a comprehensive set of experiments is presented to evaluate the proposed method’s performance, both during training and under varied test conditions. Different dataset types and sample sizes are examined to determine their impact on learning efficiency and convergence, while generalization to unseen environments is also assessed.
@@
The dataset generation and the evaluation experiments are conducted using the EV2Gym simulator (Orfanoudakis et al., 2024a), which leverages real-world data distributions, including EV arrivals, EV specifications, electricity prices, etc. This setup ensures a realistic environment where the state and action spaces accurately reflect real charging stations’ operational complexity. A scenario with 25 chargers is chosen, allowing up to 25 EVs to be connected simultaneously. In this configuration, the action vector has up to 25 variables (one per EV), while the state vector contains around 150 variables describing EV statuses, charger conditions, power transformer constraints, and broader environmental factors. Consequently, the resulting optimization problem is in the moderate-to-large scale range, reflecting the key complexities of real-world EV charging infrastructure. Each training procedure is repeated 10 times with distinct random seeds to ensure statistically robust findings. All reported rewards represent the average performance over 50 evaluation scenarios, each featuring different configurations (electricity prices, EV behavior, power limits, etc.).
@@
Offline RL algorithms, including DTs, can learn optimal policies from trajectories without needing online interaction with the environment. Consequently, the quality of the gathered training trajectories has a substantial impact on the learning process. In this work, three distinct strategies were used to generate trajectories:
  • Random Actions: Uniformly sampled actions in the range $[-1,1]$ were applied to the simulator.
  • Business-as-Usual (BaU): A Round Robin charging policy commonly employed by CPOs, which sequentially allocates charging power among EVs to balance fairness and efficiency.
  • Optimal Policies: Optimal solutions derived from solving offline the mathematical problem described in Section 3.2 for randomly generated scenarios.
Each trajectory consists of 300 state-action-reward-action mask tuples, with each timestep representing a 15-minute interval, resulting in a total of three simulated days. This combination of random, typical, and expert data provides a comprehensive basis for evaluating how GNN-DT learns from diverse offline trajectories.
Table 1: Comparison of maximum episode rewards (×$\times 10^{5}$) achieved by baseline methods and GNN-DT across various datasets and context lengths ($K$). Rewards highlighted in bold indicate the highest value within each dataset and $K$ category.
@@
5.1 Training Performance
Fig. 2 compares the proposed GNN-DT against two baselines, classic DT [^4] https://github.com/kzl/decision-transformer and Q-DT [^5] https://github.com/charleshsc/QT, which both rely on flattened state representations due to their inability to directly process graph-structured data. In these baseline methods, empty chargers and unavailable actions are replaced by zeros, so the action vector is always the same size. Several well-known online RL algorithms from the Stable-Baselines 3 framework (Raffin et al., 2021) (SAC, DDPG, TD3, TRPO, PPO, and TQC) are also included to provide a performance benchmark for complex optimization tasks featuring both dense and sparse rewards. The offline RL algorithms (DT, Q-DT, and GNN-DT) are trained on three different datasets (Optimal, Random, and BaU), each comprising 10.000 trajectories. A red dotted line marks the oracle reward, which represents the experimental maximum achievable reward obtained by completing a full simulation without uncertainty. This oracle reward serves as an upper bound and helps contextualize the relative performance of each method.
In Figs. 2.a-c, the offline DT-based approaches use a context length $K=10$ and learn from pre-collected trajectories. As expected, the Optimal dataset provides the highest-quality information, enabling GNN-DT to converge rapidly toward near-oracle performance, while classic DT and Q-DT lag far behind, showcasing GNN-DTs improved sampling efficiency. With the Random dataset, the limited quality of data leads all methods to plateau at lower reward values, although GNN-DT still surpasses the other baselines. An intriguing behavior is observed in the BaU dataset, where classic DT initially experiences a substantial drop but later recovers to a final reward level exceeding that of Q-DT and GNN-DT. In contrast, the online RL algorithms displayed in Fig. 2.d struggle to achieve comparable improvements, suggesting that pure online exploration is insufficient for solving this complex EV charging optimization problem with sparse rewards. For completeness, Appendix B.2 contains the training curves for all algorithm-dataset-context length configurations used.
@@ 5.1 Training Performance
Table 2: Maximum reward of GNN-DT trained on merged Optimal and Random datasets for $K\!=\!2$ and $K\!=\!10$. Notice that performance improves even though the average training dataset rewards are substantially lower, showing the importance of using diverse datasets in training DTs. Rewards highlighted in bold represent the highest values within each $K$ category.
@@
The previous section highlighted that the quality of trajectories in the training dataset is the most influential factor for achieving high performance. In this section, we explore whether creating new datasets by mixing existing ones can further improve performance. Initially, the Optimal and Random datasets are combined in different proportions, as summarized in Table 2. A noteworthy result is that supplementing the Optimal dataset with theoretically less useful (Random) trajectories consistently boosts performance. In particular, GNN-DT with $K=10$, trained on a mix of 250 Optimal and 750 Random trajectories, achieves near-oracle results, deviating by only $-0.001\times 10^{5}$ from the optimal reward. A similar trend emerges when blending BaU and Random datasets (Table 3), although the performance gains are not as significant. Overall, these findings indicate that carefully integrating high- and lower-quality data can enhance policy learning beyond what purely Optimal or purely Random datasets can provide.
Table 3: Maximum reward of GNN-DT trained on merged BaU and Random datasets for $K=2$ and $K=10$.
Dataset Total Traj. Avg. Dataset Reward GNN-DT Reward (×$\times 10^{5}$)
K=2 k=10
Random (Rnd.) 100% $1000$ $-2.37$$0.39$ $-0.863$ $-1.187$
BaU 25% + Rnd. 75% $1000$ $-1.93$$0.80$ $-0.578$ $-0.461$
BaU 50% + Rnd. 50% $1000$ $-1.51$$0.87$ $-0.665$ $\mathbf{-0.447}$
BaU 75% + Rnd. 25% $1000$ $-1.09$$0.76$ $\mathbf{-0.421}$ $-0.471$
BaU 100% $1000$ $-0.01$$\textpm 0.01$ $-0.654$ $-0.572$
Impact of larger context lengths (K)
Figure 3: GNN-DT performance for larger context lengths (K).

### v2

@@ 4 GNN-based Decision Transformer
Figure 1: Overview of the GNN-DT architecture. The input sequence, comprising return-to-go, action, and state, is processed through specialized embedding modules. The action graph $\mathcal{G}^{\mathbf{a}}_{t}=(\mathcal{N}^{\mathbf{a}}_{t},\mathcal{E}^{\mathbf{a}}_{t})$, with nodes $\mathcal{N}^{\mathbf{a}}_{t}\subset\mathcal{N}_{t}$, and the state graph $\mathcal{G}_{t}=(\mathcal{N}_{t},\mathcal{E}_{t})$ are encoded using GNN-based embedders to produce embeddings of dimension $F_{L}$. These embeddings serve as inputs to a GPT-2–based causal transformer, which predicts the next action token. The predicted action token acts as a decoder, generating actions by multiplying with specific GNN state node embeddings.
@@ 4.1 Sequence Embeddings
In detail, the State Embedder consists of $L$ consecutive Graph Convolutional Network (GCN) [36] layers, which aggregate information from neighboring nodes as follows:
@@
where $\boldsymbol{x}_{t}^{(l)}\in\mathbb{R}^{N_{t}\times F_{l}}$ denotes the node embeddings at layer $l$ with $N_{t}$ number of nodes, $W^{(l)}\in\mathbb{R}^{F_{l}\times F_{l+1}}$ are trainable weights, $\sigma(\cdot)$ is a nonlinear activation (ReLU), $A_{t}$ is the adjacency matrix of the state graph $\mathcal{G}_{t}$, and $D$ is the degree matrix for normalization. After the final layer, a mean-pooling operation produces a fixed-size state embedding: $\boldsymbol{\widetilde{s}}_{t}=\frac{1}{\lvert\mathcal{N}_{t}\rvert}\sum_{n\in\mathcal{N}_{t}}\boldsymbol{x}_{n,t}^{(L)},$ where $\boldsymbol{x}_{n}^{(L)}$ is the embedding of node $n$ at the $L$-th layer. This pooling step ensures that the state embedding is invariant to the number of nodes in the graph, enabling the architecture to scale with any number of EVs or chargers. Similarly, the Action Embedder processes the action graph $\mathcal{G}^{\mathbf{a}}_{t}=(\mathcal{N}^{\mathbf{a}}_{t},\mathcal{E}^{\mathbf{a}}_{t})$ through $C$ GCN layers followed by mean pooling, producing the action embedding $\widetilde{\textbf{a}}_{t}$. All embedding vectors (states, actions, or the return-to-go value) have the same dimensions. This design leverages the dynamic and invariant nature of GCN-based embeddings, allowing the DT to handle variable-sized graphs.
@@ 4.2 Decoding Actions
Once the embedding sequence of length $K$ is constructed [^2] During inference the action ($\mathbf{a}_{t}$) and RTG ($\boldsymbol{\widehat{R}}_{t}$) of the last step $t$ are filled with zeros as they are not known., it is passed through the causal transformer GPT-2 to produce a fixed-size output vector $\boldsymbol{y}_{t}\in\mathbb{R}^{F_{L}}$ for each step. Because DT architectures inherently generate outputs of fixed dimensions, an additional mechanism is required to manage dynamic action spaces. To address this, GNN-DT implements a residual connection that merges the final GCN layer embeddings $\boldsymbol{x}_{t}^{(L)}$ with the transformer output $\boldsymbol{y}_{t}$ for every step of the sequence. Specifically, for each node $n\in\mathcal{N}_{t}^{\mathbf{a}}$, we retrieve its corresponding state embedding $\boldsymbol{x}_{n,t}^{(L)}\in\mathbb{R}^{1\times F_{L}}$ and multiply it with the transformer output token $\boldsymbol{y}_{t}\in\mathbb{R}^{1\times F_{L}}$, yielding the final action for node $n$: $\hat{\mathrm{a}}_{n,t}=\boldsymbol{y}_{t}^{\mathsf{T}}\cdot\boldsymbol{x}_{n,t}^{(L)}.$ By repeating for every step $t$ and every node $n\in\mathcal{N}^{\mathbf{a}}_{t}$ the final action vector $\hat{\mathbf{a}}_{t}$ is generated. This design allows the model to maintain a fixed-size output from the DT while dynamically adapting to any number of nodes (and hence actions). It effectively combines the high-level context learned by the transformer with the node-specific state information captured by the GNN, enabling robust, scalable decision-making even as the graph structure changes.
4.3 Training: Action Masking and Loss Function
The proposed GNN-DT model is trained via supervised learning using an offline trajectory dataset [13], similarly to offline RL. Specifically, the GPT-2 model is initialized with its default pre-trained weights, which are subsequently fine-tuned end-to-end for the EV charging optimization task. In GNN-DT, the learning of infeasible actions, such as charging an unavailable EV, is avoided through action masking. At each time step $t$, a mask vector $\mathbf{m}_{t}$, which has the same dimension as $\mathbf{a}_{t}$, is generated with zeros marking invalid actions and ones marking valid actions. For example, an action is invalid when the $a_{i,t}\neq 0$ and no EV is connected at charger $i$. The mean squared error between the predicted actions $\widehat{\mathbf{a}}_{t}$ and ground-truth actions $\mathbf{a}_{t}$ from expert or offline trajectories is employed as the loss function. For a window of length $K$ ending at time $t$, training loss is defined as:
$\mathcal{L}=\frac{1}{K}\sum\nolimits_{\tau=t-K}^{t}\bigl{\lVert}(\widehat{\mathbf{a}}_{\tau}-\mathbf{a}_{\tau})\circ\mathbf{m}_{\tau}\bigr{\rVert}^{2}.$ (6)
By incorporating the mask into the loss calculation (elementwise multiplication), a focus solely on valid actions is enforced, thereby preserving meaningful gradient updates.
@@ 5 Experiments
In this section, a comprehensive set of experiments is presented to evaluate the proposed method’s performance, both during training and under varied test conditions. Different dataset types and sample sizes are examined to determine their impact on learning efficiency and convergence.
@@
The dataset generation and the evaluation experiments are conducted using the EV2Gym simulator [17], which leverages real-world data distributions, including EV arrivals, EV specifications, electricity prices, etc. This setup ensures a realistic environment where the state and action spaces accurately reflect real charging stations’ operational complexity. A scenario with 25 chargers is chosen, allowing up to 25 EVs to be connected simultaneously. In this configuration, the action vector has up to 25 variables (one per EV), while the state vector contains around 150 variables describing EV statuses, charger conditions, power transformer constraints, and broader environmental factors. Consequently, the resulting optimization problem is in the moderate-to-large scale range, reflecting the key complexities of real-world EV charging. Each training procedure is repeated 10 times with distinct random seeds to ensure statistically robust findings. All reported rewards represent the average performance over 50 evaluation scenarios, each featuring different configurations (electricity prices, EV behavior, power limits, etc.). Training was carried out on an NVIDIA A10 GPU paired with 11 CPU cores and 80 GB of RAM, using the AdamW optimizer and a LambdaLR scheduler. Baseline RL agents converged in 2–5 hours, while the proposed GNN-DT required up to 24 hours of training. Default hyperparameters were used for all baseline RL methods; Appendix B.1 lists the full set of hyperparameters employed to train the DTs.
@@
Offline RL algorithms, including DTs, can learn policies from trajectories without needing online interaction with the environment. Consequently, the quality of the gathered training trajectories has a substantial impact on the learning process. In this work, three distinct strategies were used to generate trajectories: a) Random: Uniformly sampled actions in the range $[-1,1]$ were applied to the simulator. b) Business-as-Usual (BaU): A Round Robin heuristic charging policy commonly employed by CPOs, which sequentially allocates charging power among EVs to balance fairness and efficiency. c) Optimal: Expert solutions derived from solving offline the MIP as described in Section 3.2 for multiple scenarios. Each trajectory consists of 300 state-action-reward-action mask tuples, with each timestep representing a 15-minute interval, resulting in a total of three simulated days. This combination of random, typical, and expert data provides a comprehensive basis for evaluating how GNN-DT learns from diverse offline trajectories.
5.1 Training Performance
Fig. 2 compares the proposed GNN-DT against multiple baselines, including the classic DT [13] and Q-DT [15] , which both rely on flattened state representations due to their inability to directly process graph-structured data. In these baseline methods, empty chargers and unavailable actions are replaced by zeros, so the action vector is always the same size. Several well-known online RL algorithms from the Stable-Baselines-3 [37] framework (SAC, DDPG, TD3, TRPO, PPO, and TQC) and offline RL algorithms from D3RLPY [38] (IQL, CQL, and BC) are also included. The offline RL algorithms (IQL, CQL, BC, DT, Q-DT, and GNN-DT) are trained on three datasets (Optimal, Random, and BaU), each comprising 10.000 trajectories. A red dotted line marks the optimal reward, which represents the experimental maximum achievable reward obtained by solving the deterministic MIP knowing the future (Oracle) defined in Appendix A.1. This oracle reward serves as an upper bound and helps contextualize the relative performance of each method. It is important to note that the EV smart charging problem requires real-time (1-5 min intervals) optimization solutions for large-scale, highly stochastic scenarios, where metaheuristic algorithms, stochastic optimization, and model predictive control methods fail due to computational constraints. By contrast, once trained (over 2–24 hours), RL agents can deliver real-time charging schedules on an ordinary computer in milliseconds.
Figure 2: Training performance comparison for online and offline RL algorithms.
In Figs. 2.a-c, the DT-based approaches use a context length $K=10$. As expected, the Optimal dataset provides the highest-quality information, enabling GNN-DT to converge rapidly toward near-oracle performance, while classic DT, Q-DT, and the other offline RL algorithms lag far behind, showcasing GNN-DT’s improved sample efficiency. With the Random dataset, the limited quality of data leads all methods to plateau at lower reward values, although GNN-DT still surpasses the other baselines. An intriguing behavior is observed with the BaU dataset, where classic DT, BC, and IQL converge at rewards exceeding that of GNN-DT. In contrast, the online RL algorithms displayed in Fig. 2.d struggle to achieve comparable improvements, suggesting that pure online exploration is insufficient for solving this complex EV charging optimization problem with sparse rewards. For completeness, Appendices B.2 and B.3 contain the training curves for all algorithm-dataset-context length configurations used. In the rest of this section, we omit the online and offline RL baselines, as their performance is substantially inferior to that of DT, Q-DT, and GNN-DT.
Dataset Impact
Table 1: Comparison of maximum episode rewards (×$\times 10^{5}$) for baselines and GNN-DT across datasets and context lengths ($K$). Bold indicates the highest value within each dataset and $K$ category.
@@ 5.1 Training Performance
The previous section highlighted that the quality of trajectories in the training dataset is the most influential factor for achieving high performance. In this section, we explore whether creating new datasets by mixing existing ones can further improve performance. The Optimal and Random datasets are combined in different proportions, as summarized in Table 2. A noteworthy result is that supplementing the Optimal dataset with theoretically less useful (Random) trajectories consistently boosts performance. In particular, GNN-DT with $K=10$, trained on a mix of 250 Optimal and 750 Random trajectories, achieves near-oracle results, deviating by only $-0.001\times 10^{5}$ from the optimal reward. A similar trend emerges when blending BaU and Random datasets (Table 8 in Appendix B.4), although the performance gains are not as significant. Overall, these findings indicate that carefully integrating high- and lower-quality data can enhance policy learning beyond what purely Optimal or purely Random datasets can provide.
Table 2: Maximum reward of GNN-DT trained on merged Optimal and Random datasets for $K=2$ and $K=10$. Highest rewards per $K$ are highlighted with bold.
@@
Component Ablation Study
To better understand the contribution of each architectural component, we conduct an ablation study that systematically removes or replaces elements of our model and then trains the model on the Optimal and Mixed (Opt.25%+Rand.75%). The results in Table 3 reveal that neither a plain DT nor a DT augmented solely with a state-GNN submodule achieves competitive performance. Notably, adding the residual connection atop the state-GNN leads to a significant improvement, from $-0.77\times 10^{5}$ to $-0.16\times 10^{5}$ on the Mixed dataset, demonstrating its importance for effective credit assignment over dynamic inputs. Removing action masking or replacing the GCN module with a Graph Attention Network (GAT) [39] similarly degrades performance, indicating that each component provides distinct and complementary benefits. Ultimately, only the full GNN-DT architecture achieves strong performance across both the Mixed and Optimal datasets.
Table 3: Average reward trained over 5 runs with different seeds for Optimal and Mixed datasets.
{tblr}
cells = c, hline1,8 = -0.08em, hline2 = -0.05em, DT & State GNN Action GNN Res. Con. Action Mask Optimal (×$\times 10^{5}$) Mixed (×$\times 10^{5}$)
✓ ✗ ✗ ✗ ✗ $-0.69\pm 0.03$ $-0.95\pm 0.39$
✓ GCN ✗ ✗ ✗ $-0.71\pm 0.02$ $-0.77\pm 0.17$
✓ GCN ✗ ✓ ✗ $-0.18\pm 0.03$ $-0.16\pm 0.07$
✓ GCN GCN ✓ ✗ $-0.11\pm 0.03$ $-0.12\pm 0.04$
✓ GAT GAT ✓ ✓ $-0.14\pm 0.07$ $-0.15\pm 0.06$
✓ GCN GCN ✓ ✓ $\mathbf{-0.09\pm 0.02}$ $\mathbf{-0.10\pm 0.04}$
5.2 Illustrative Example of EV Charging
With the models trained, we proceed to compare the behavior of the best baseline models trained (DT, Q-DT, GNN-DT) against the heuristic BaU algorithm in an EV charging scenario. Fig. 3a presents the SoC progress for three EVs connected one after the other to a single charger throughout the simulation, while Fig. 3b shows the actions of all chargers taken by each algorithm.